# COCOLA: Coherence-Oriented Contrastive Learning of Musical Audio Representations

### v2

@@ 4.3 Training Details
All COCOLA models are trained on an NVIDIA RTX 4070 Super with 12GB of VRAM. Each training batch contains 32 audio chunks of 5s (16kHz). We set the maximum window overlap ratio $r=50\%$ and train with the Adam optimizer [68] with a $10^{-3}$ learning rate. We add Gaussian noise to positive samples as a data augmentation method, with $\sigma=10^{-3}$. We tried training with cosine similarity [22] but we reported (Figure 4) lower performance, corroborating [33].
CompoNet models are also fine-tuned on NVIDIA GPUs (RTX 4080 16GB on MUSDB18-HQ; A10G, 24GB on Slakh2100). The training batches have size 2 (MUSDB18-HQ) and 5 (Slakh2100) and contain windows of 10.24s (16kHz). We train with Adam [68] with a $10^{-4}$ learning rate. We do not use the empty prompt for classifier-free guidance [69] during fine-tuning.
5 Experiments
@@
We employ four COCOLA encoder models in our experiments : “COCOLA MoisesDB”, “COCOLA Slakh2100”, “COCOLA CocoChorales” and “COCOLA All”. The first three are trained on the homonym datasets, while the last one is trained on all three combined. For the “COCOLA CocoChorales” we use all ensables while on “COCOLA All” we use only the Random ensamble for a more balanced partitioning, with respect to the other datasets. MUSDB18-HQ, being the smallest dataset, is used as a held-out test dataset for studying generalization.
5.1 Coherent Sub-Mix Classification
@@
We cross-test the performance of all COCOLA models, performing classification of coherent pairs on the test split of our datasets. More specifically, given an encoder $f_{\theta}$, we iterate a test set, collecting at each step a batch of $K$ windows $\mathbf{x}^{1},\dots,\mathbf{x}^{K}$. Following the steps in Section 3.1 we compute all similarities $\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2})$ for $k,j\in[K]$. We define the accuracy over a batch as:
@@
where $\mathds{1}$ is the indicator function. We obtain the final accuracy averaging over all batches in the dataset. For our evaluation we use $K=2$ and depict in Table 2 the results across all combinations of models and test datasets. While both “COCOLA MoisesDB” and “COCOLA Slakh2100” perform only slightly better than a random choice, “COCOLA CocoChorales” features improved performance. Finally, combining the three dataset, we obtain an accuracy of over 90% on all datasets, showcasing generalization with 90.43% on the held-out MUSDB18-HQ.
5.2 Accompaniment Generation Evaluation
For the accompaniment generation evaluation, we compare the MSDM model [11] with the proposed CompoNet (Section 3.3). We train CompoNet on MUSDB18-HQ and Slakh2100 (restricted to Bass, Drums, Guitar and Piano stems at test time). We also consider a Random baseline, where, for a given input, we output a random sub-mix from a different test track. We generate 200 chunks for both datasets and models, conditioning on random stem subsets of test tracks and querying a subset of the complementary. The chunks are $\sim$6s / 10.24s long on MUSDB18-HQ / Slakh2100. Given that MSDM tends to generate silence, we sample 12 candidate tracks for each generated track, selecting the one with the highest $L^{2}$ norm. We compare the COCOLA score in Eq. (4) with the FAD [16, 70] metric (interpreted as a sub-FAD [15, 11]) computed with CLAP [9], EnCodec [71], and VGGish [72] backbones.
We showcase the results in Table 3. With the FAD metrics, the model assigns the best score to the Random baseline. This behavior can be explained by considering that the Random outputs are real data, and the FAD evaluates well the perceptual quality. At the same time, it fails to assess the coherence between the tracks, and tends to score MSDM better. With the “COCOLA All” score, Random is the lowest, while CompoNet scores best. CompoNet has the highest value on “COCOLA CocoChorales” which, however, scores MSDM lower than Random. We remember that “COCOLA All” achieves a higher accuracy in Table 2, resulting in more reliability. On Ground Truth, we compute upper-bounds of the COCOLA score using real positive pairs. CompoNet better approaches these bounds.
6 Conclusion
In this paper we proposed COCOLA, a contrastive encoder for recognizing the coherence between musical stems. At the same time, we have introduced CompoNet, a novel compositional model for music based on ControlNet that can solve a wide range of tasks simultaneously. Finally, we have evaluated CompoNet using COCOLA, proposing a new way of assessing accompaniment generation.
We plan to improve the quality of COCOLA by training on additional stem-level datasets [73] or using data obtained by pre-separating [15] larger realistic music datasets [74]. In future work, we would also like to explore inference-side methods that can guide the diffusion process using COCOLA as a likelihood function, offering an alternative (or additional) loss for the GMSDI method [46].
7 Acknowledgments
The authors were partially supported by the ERC grant no. 802554 (SPECGEO), PRIN 2020 project no. 2020TA3K9N (LEGO.AI), PRIN 2022 project no. 2022AL45R2 (EYE-FI.AI, CUP H53D2300350-0001), PNRR MUR project no. PE0000013-FAIR, and AWS Cloud Credit for Research program.

### v3

@@ IV-C Training Details
All COCOLA models are trained on an NVIDIA RTX 4070 Super with 12GB of VRAM. Each training batch contains 32 audio chunks of 5s (16kHz). We set the maximum window overlap ratio $r=50\%$ and train with the Adam optimizer with a $10^{-3}$ learning rate. As a data augmentation method, we add Gaussian noise to positive samples, with $\sigma=10^{-3}$.
During the training with HPS, masking is randomly applied to either one of the two channels or none. This approach allows the model to learn representations from either component independently.
V Experiments
@@
In our experiments, we trained four COCOLA encoder models without HPS: “COCOLA MoisesDB”, “COCOLA Slakh2100”, “COCOLA CocoChorales” and “COCOLA All”. The first three are trained on the homonym datasets, while the last one is trained on all three combined. For the “COCOLA CocoChorales” we use all ensembles, while on “COCOLA All” we use only the Random ensemble for a more balanced partitioning with respect to the other datasets. Additionally, we train a “COCOLA HPS All” model employing HPS with the same data as “COCOLA ALL”. MUSDB18-HQ, being the smallest dataset, is used as a held-out test dataset for studying generalization.
V-A Coherent Sub-Mix Classification
@@
We cross-test the performance of COCOLA models trained without HPS by classifying coherent pairs on the test splits of our datasets. More specifically, given an encoder $f_{\theta}$, we iterate a test set, collecting at each step a batch of $K$ windows $\mathbf{x}^{1},\dots,\mathbf{x}^{K}$. Following the steps in Section III-A we compute all similarities $\text{sim}(\mathbf{h}^{k}_{1},\mathbf{h}^{j}_{2})$ for $k,j\in[K]$. We define the accuracy over a batch as:
@@
where $\mathds{1}$ is the indicator function. We compute the final accuracy by averaging across all batches. Using $K=2$, Table I shows results for various model and test dataset combinations. While ”COCOLA MoisesDB” and ”COCOLA Slakh2100” perform slightly better than random, ”COCOLA Coco-Chorales” shows improved performance. Combining all three datasets achieves over 90% accuracy.
Adding HPS to “COCOLA All”, all metrics improve except on CocoChorales (which is already high), reaching 93.87% on the held-out MUSDB18-HQ and showcasing the usefulness of the factorized representation. We also investigate the performance of “COCOLA HPS All” increasing $K$ in Table II.
V-B Accompaniment Generation Evaluation
For the music accompaniment generation evaluation, we compare GMSDI [32], Diff-A-Riff [28], and Stable Audio ControlNet (SA ControlNet). The latter is an implementation of a DiT-based [41] ControlNet [42] for Stable Audio Open [4], which we release [^3] https://github.com/EmilianPostolache/stable-audio-controlnet, given the lack of state-of-the-art open-source models for music accompaniment generation (only MSDM [9] is available, but it is trained on Slakh2100 [36] not being applicable in realistic settings).
GMSDI is trained on MTG-Jamendo [43], Diff-A-Riff on a proprietary dataset and we finetune SA ControlNet on MoisesDB. We also consider a Random baseline, where, for a given input, we output a random stem from a different test chunk. We evaluate the models both on the MoisesDB and MUSDB18-HQ test splits.
We sample 168 chunks ($\sim$47.55s) from the MoisesDB test split and 236 chunks of the same length from the MUSDB18-HQ test split, using random stem subsets as input and querying complementary stems, excluding vocals from the output and the ”Other” class in MUSDB18-HQ. SA ControlNet processes the full chunk, while GMSDI and Diff-A-Riff use shorter initial sub-chunks ($\sim$23.78s and $\sim$9.58s, respectively) due to reduced context windows. Outputs are cut to the initial $\sim$9.58s.
We compare the COCOLA Score (based on ”COCOLA HPS All”) with the FAD [14, 44] metric (interpreted as a sub-FAD [13, 9]), computed using CLAP [7], EnCodec [45], and VGGish. Results in Table III show that FAD tends to favor the Random baseline, as it evaluates perceptual quality but not track coherence, which is better captured by COCOLA. On Ground Truth, we compute COCOLA upper bounds with real positive pairs. Diff-A-Riff performs best on MoisesDB, and SA ControlNet on MUSDB18-HQ.
V-C MOS vs COCOLA Score
To evaluate COCOLA’s effectiveness, we conducted a subjective test with 41 participants who rated (0-4) the coherence of (pitch) Shifted and (time) Stretched chunks. We also used Random, Ground Truth, Diff-A-Riff and SA ControlNet tracks from the previous experiment. We average the ratings (MOS) for each chunk. Figure 3 shows the correlation plot between MOS and COCOLA Score (using ”COCOLA HPS All”). The Pearson correlation coefficient is $r=$ 0.54 $(p<0.01)$ indicating a good correlation with high statistical signficance.
VI Conclusion
In this paper, we proposed COCOLA, a contrastive encoder for recognizing the coherence between musical stems. Both classification experiments and human scores demonstrate the efficacy of the method. We used COCOLA for benchmarking music accompaniment models, proposing a new evaluation metric for the task. We plan to improve the quality of COCOLA by training on additional stem-level datasets [46] or on pre-separated [13] large scale datasets [43].